OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio three colon two. If the library has three hundred sixty fiction books, how many nonfiction books does it have? seventy-two one hundred twenty one hundred eighty two hundred forty

OpenStudy (anonymous):

3 parts of fiction and 2 parts of non fiction 3x+2x=360

OpenStudy (anonymous):

similar to prev one

OpenStudy (anonymous):

fiction/ non fiction = 3/2 = 3x/2x 3x=360, x = 120 non fiction = 2x = 240